OpenStudy (anonymous):

Find the derivative dy/dx = y' using Implicit Differentiation: x^2 - y^2 = 16

OpenStudy (turingtest):

why don't you try to take the derivative and show me what you get so I can tell you where you go wrong

OpenStudy (anonymous):

Is this the answer: 2x-2yy'=0

OpenStudy (turingtest):

yes, great! but we are not done because we have not found dy/dx so now solve for y'

OpenStudy (anonymous):

yeah that's the part where I don't know.

OpenStudy (turingtest):

it's just algebra, perhaps the notation is confusing you what if it was 2x-2yd=0 and I asked you to solve for d ? could you do that?

OpenStudy (anonymous):

y' = 2x/2y? So these problems have two steps?

OpenStudy (amistre64):

what do you mean by "two steps"?

OpenStudy (turingtest):

or three usually now we need to plug in for y

OpenStudy (turingtest):

...not necessarily, but often they want you to get everything in terms of x -amistre: I think he is counting 1)taking derivative 2)solving for y'

OpenStudy (amistre64):

subbing in for y would have made this more explicit to begin with if you could solve for y ;)

OpenStudy (turingtest):

...but it does say to differentiate implicitly -so the next step would be 3)sub in for y\[x^2-y^2=16\implies y=\sqrt{x^2-16}\]

OpenStudy (anonymous):

so put the y'=2x/2y into where?

OpenStudy (anonymous):

i dont no

OpenStudy (turingtest):

first of all, you can't neglect basic algebra we found that\[y'=\frac{\cancel2^1x}{\cancel2y}=\frac xy\]do try to remember thigns like the fact that the 2's canceland we have from above that\[x^2-y^2=16\implies y=\sqrt{x^2-16}\]so plugging that expression for y' we get\[y'=\frac xy={x\over\sqrt{x^2-16}}\]